5. Suppose a consumer has income of 200 birr per month and he wants to spend


all of his income on two goods- X and Y whose prices are birr 4 and birr 5


respectively. Based on this information answer the following questions:


a. Express the budget line of the consumer both algebraically and



diagrammatically.


b. Compute the equation of the budget line.



c. Determine the slope of the budget line and interpret the result



e. What will happen to the original budget line?

1
Expert's answer
2022-01-20T16:15:37-0500

a)Budget line expressed algebraically:

"M=p_xQ_x+p_yQ_y"

Diagrammatically



b)The equation of the budget line

"M=PxQx+PyQy\\



Qy=\frac{M}{p_y}-frac{P_x}{P_y} ×Q_x"


M=income

"P_y" =price of y

"P_x" =price of x

"Q_y=\frac{200}{5}-\frac{5}{5x}"

"Q_y=40-x"


c)The slope of the budget line

slope="-\frac{P_x}{P_y}"

"=\frac{0-40}{40-0}\\

=\frac{-40}{40}\\

=-1"

e)What will happen to the original budget line

if money income doubles the original budget line will not move .Exactly the same quantities can be purchased as before

if the price levels will increase then the budget line will move inwards

if the price of y doubles the line would move outward on the point where the budget line crosses the horizontal axis